OpenStudy (anonymous):

I will fan and medal for help! What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4 (blank, blank)

OpenStudy (anonymous):

it says on Algebra.com x= 15/-5 for the answer. I copied your question and found that.

OpenStudy (anonymous):

are you sure...

OpenStudy (anonymous):

do you know @dolphin1998 ?

OpenStudy (anonymous):

I am sure I will show you what they got.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

SINCE Y=-X+4 THEN WE SUBSTITUTE (-X+4) FIR Y IN THE OTHER EQUATION WE GET -X+4=2X/3-1 -X-2X/3=-1-4 (-3X-2X)/3=-5 -5X/3=-5 NOW CROSS MULTIPLY -5X=-15 X=-15/-5 X=3 ANSWER. NOW SUBSTITUTE 3 FOR X & SOLVE FOR Y Y=-3+4 Y=1 ANSWER. PROOF 1=2/3*3-1 1=2-1 1=1 GRAPHS FOLLOW

OpenStudy (anonymous):

so basically your answer is x=3 and y=1 for the same problem.

OpenStudy (dolphin1998):

IvyCH is right i looked it up too

OpenStudy (anonymous):

so in the blanks its 3 and 1?

OpenStudy (anonymous):

XD man im so clue-less XD

OpenStudy (anonymous):

yes, it should be since it says X= 3 and Y= 1 for your answers

OpenStudy (anonymous):

can you help with another...?

OpenStudy (anonymous):

what's next.

OpenStudy (anonymous):

Alex graphed the system of equations and stated that the solution is (5, 2). Check his solution algebraically. https://static.k12.com/calms_media/media/1513500_1514000/1513724/1/e43309f64710a063e8b7127f2236dbf13670ffda/MS_IMC-1400804-100531.jpg Which statement is true? A. The graph for y=-3/5x+5 is incorrect. B. The graph for y=2/5x is incorrect. C. (5, 2) is the solution to the system of equations. D. (2, 5) is the solution to the system of equations.

OpenStudy (anonymous):

do any of you guys know this one...? XDDD I see many of you are looking at this question...

OpenStudy (anonymous):

I am looking it up for you.

OpenStudy (anonymous):

someone else had the same question and someone answered it saying D. (2,5) is the solution to the system of equations.

OpenStudy (anonymous):

...alright...thanks

OpenStudy (anonymous):

welcome, I am still looking to see alot of people have this question.

OpenStudy (anonymous):

everyone's saying D

OpenStudy (anonymous):

ya, its most likely d...thanks again